OpenStudy (anonymous):

help finishing a derivative problem... f(x)= (sqrt(x)+52)(sqrt(x)-51) find deriv. I calculated this to the following:

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

i would take the derivative of \[\frac{x+52}{x-51}\] then replace \(x\) by \(\sqrt{x}\) and multiply the result by \(\frac{1}{2\sqrt{x}}\)

OpenStudy (anonymous):